Solve. Write empty set of no solution exists
| x - 1/2 | = 3/2

Answer by stanbon(75887)

| x - 1/2 | = 3/2
x-1/2 = 3/2 OR x-1/2 = -3/2
-----
x = 2 OR x = -1
----------------
Cheers,
Stan H.
-=---------